Sol Summary – January 29th

Sol: 8
Summary Title: Extending the search
Author’s name: Andrew Wheeler
Mission Status: Nominal
Sol Activity Summary: Day dawned clear and cold (-5C) and augured well for the crew’s plans. Rob finally had both clear skies and a fully operational telescope to observe the sun. The resultant images are being processed and the astronomer’s report will have the details. The illnesses of yesterday are where they belong – yesterday. A localised morning EVA around the hab further tested the robot rover’s mobility and its ability to gather spectra whilst being remotely operated. After lunch, an EVA to the furthest north reached by the rovers at Tharsis Montes had a team searching for evaporites, concretions and petrified wood. Serendipitously, as the team was looking for rock markers on the top of Hab ridge, signs of live were discovered. Endoliths were observed beneath the surface of quartz cobbles. Three samples were selected for spectroscopic analysis and comparison to the vegetation spectra from the increasing MDRS-derived spectral library. The evening meal provided an opportunity to clear the shelves of opened foodstuffs with a simple meat and seven vegetables spread.
Look Ahead Plan: As the forecast predicts more clear and slightly warmer days, we intend to push as far north as the rovers’ limit will allow in our attempts to reach the white moon gypsum deposit to collect samples. A second EVA will examine anhydrite yardangs in an environmental degradation study.
Anomalies in work: Nothing to report.
Weather: Clear morning at -5 Celsius rising to mid to high teens during the day. No clouds and no precipitation. Humidity continuously dropping from near 60% to 30%. The wind was faint to gentle during the day.
Crew Physical Status: Nothing to report.
EVA: EVA 13 with Steve and Scott with robot rover trials at the Hab.
EVA 14 with Andrew, Clare and Alex to Tharsis Montes in search of evaporites, concretions and petrified wood.
Reports to be filed: Sol Summary, Commander’s Report, Greenhab Report, Operations Report (including an HSO report), EVA report, EVA Request(s), Astronomer report.
Support Requested: Nothing to request.

Copyright © The Mars Society. All rights reserved. | Main Site